madison and alexa have a combined age of 32. madison is older, and the difference in their ages is 4 years. how old is each person?
Answer by mananth(16946)

Madison's age be x and alexa's age be y
madison and alexa have a combined age of 32.
x+y =32

Madison is older, and the difference in their ages is 4 years.
x-y=4
x=y+4
substitute x in x+y=32
y+4+y =32
2y+4=32
2y=28
y = 14
Alexa is 14 years
and Madison is 4 years older. 18 years